1. Ordered: ASA gr XV PO QID

Available: ASA 325 mg/tablet

Give:_________ How many times per day? ________

2. Ordered: Tetracycline 0.5 gm PO QID

Available: Tetracycline 500 mg/capsule

Give:_________

3. Ordered: Atarax 25 mg PO Q AM

Available: Atarax 10 mg/5 ml

Give:_________

4. 3 T (Tablespoon ) = ____ tsp = ----- ml

5. Ordered: Benadryl 50 mg PO HS

Available: Benadryl 25/tablet

Give:_________
Answer #1

1. Ordered: ASA gr XV PO QID means 4 times a day orally

gr. XV = 1000mg

Available tablet= 325 mg per tablet

hence required tablets = 1000/325= 3.07 tablets 4 times a day

means almost 3 tablets qid

2.Ordered Tetracycline= 0.5 gm PO QID

qid= means 4 times a day, po= orally

1gm= 1000mg

hence 0.5 gm=500mg

Available: Tetracycline 500 mg/capsule

hence required capsules= 500/500= 1 capsule

3. Ordered: Atarax= 25 mg PO Q AM Means every day before noon by oral route

Available: Atarax 10 mg/5 ml

means 2mg/ml

Required dose in ml= 25/2 ml= 12.5 ml

4. 3 T (Tablespoon ) = 9 tsp = 45 ml

5. Ordered: Benadryl = 50 mg PO HS, orally at bedtime / half strength

Benadryl 25/tablet

Hence required tablets = 50/25= 2 tablets